OpenStudy (anonymous):

evaluate the polynomial for x=4 -5x+4

OpenStudy (yrelhan4):

just put 4 in the expression and solve..

OpenStudy (anonymous):

\[-5x+4=-5\times4+4=-20+4=-16\]

OpenStudy (yrelhan4):

-5(4) + 4= -20 + 4 = -16

OpenStudy (anonymous):

thanks